# Learning Multiplex Embeddings on Text-rich Networks with One Text Encoder

### v1

@@ 5.4 Direct Inference with an Evident Source Relation
We conduct experiments on venue recommendation (Venue-Rec) and author identification (Author-Idf) in academic networks and brand prediction (Brand-Pred) in e-commerce networks to verify the effectiveness of Metern’s direct inference ability with an evident source relation. All three tasks are formulated as matching tasks where we use PREC@1 as the metric. For Venue-Rec, Author-Idf, and Brand-Pred, we use the sv, sa, and cob representation, respectively.
(a) Citation
(b) Rec
Figure 3: The learned source relation weights for citation prediction and paper recommendation on Geology network.
The results are shown in Table 3. From the results, we can find that: Metern outperforms all the baseline methods significantly and consistently on all the tasks (except Venue-Rec on Mathematics), which demonstrates that the multiplex embeddings generated by Metern are of good quality and can be directly applied to downstream tasks. Detailed information about the tasks and experimental settings can be found in Appendix A.5.2.
Table 3: Direct inference with an evident source relation (no task-specific training) on academic networks and e-commerce networks.
@@
5.5 Learning to Select Source Relations
We perform experiments on paper recommendation, paper classification, citation prediction, and year prediction on academic networks, and item classification and price prediction on e-commerce networks to test if Metern can learn to select source relations. We use PREC@1 as the metric for paper recommendation, Macro-F1 as the metric for paper classification and item classification, and RMSE as the metric for citation prediction, year prediction, and price prediction.
Table 4: Learn to select source relations on academic networks. ($\uparrow$) means the greater the score is, the better the model is, and ($\downarrow$) is on the contrary.
@@ 5.5 Learning to Select Source Relations
Table 5: Learn to select source relations on e-commerce networks.
@@
The results on academic networks and e-commerce networks are shown in Table 4 and Table 5, respectively. We also show the learned source relation weight for citation prediction and paper recommendation in Figure 3. From the results, we can find that: 1) Metern outperforms all the baseline methods significantly and consistently on all the tasks (except citation prediction on Geology and item classification on Sports). 2) Metern can learn to select different relations for different downstream tasks, i.e.,same-author” and “same-venue” for citation prediction (author and venue can largely determine the citation of the paper), and “cited-by” and “co-cited-by” for paper recommendation (paper in “cited-by” or “co-cited-by” relations are more likely to be beneficial to researchers together). Detailed information about the tasks and experimental settings can be found in Appendix A.5.3. More results analysis can be found in Appendix A.7.
5.6 Efficiency Analysis
@@
Attribute Vanilla FT MTDNN Metern
@@
We run experiments on Geology to study the time complexity and memory complexity of Metern, comparing with Vanilla FT and MTDNN. All compared models are trained for 40 epochs on four Nvidia A6000 GPU devices with a total training batch size set as 512. We show the result in Table 6. From the result, we can find that: the time complexity and memory complexity of training Metern is on par with training Vanilla FT and MTDNN.
@@ 5.7 Multiplex Embedding Visualization
We visualize the multiplex node embeddings ${\bm{h}}_{v|r}$ learned by Metern with t-SNE Van der Maaten and Hinton (2008). The results on Geology network are shown in Figure 4. We randomly select one center node from the network, together with its neighboring nodes of the “cited-by” (cb), “same-author” (sa), and “same-venue” (sv) relations. Then, we utilize Metern to encode all the nodes (including the center node and all neighboring nodes) in different relation representation spaces (i.e., to obtain ${\bm{h}}_{v|r_{cb}}$, ${\bm{h}}_{v|r_{sa}}$, and ${\bm{h}}_{v|r_{sv}}$). In the figure, neighboring nodes of different relations are marked in different colors and shapes. From Figure 4, we can find that, in the output embedding space corresponding to one relation (e.g., the ${\bm{h}}_{v|r_{cb}}$ latent space corresponding to “cited-by” in Figure 4(a)), the center node (the blue circle)’s neighboring nodes of the corresponding relation (e.g., “cited-by” relation neighbors, marked as orange rectangles) are close to the center node. This demonstrates that Metern can capture the semantics of different relations and represent them in different latent subspaces.
@@
Figure 4: Multiplex embedding visualization on Geology network. cb, sa, and sv represent the “cited-by”, “same-author”, and “same-venue” relation respectively.
@@ 5.8 Analysis of Relation Weight $w_{r}$
In Metern, different relations sometimes have different learning speeds, and $w_{r}$ in Eq.(6) can be tuned to balance the learning speed of different relations. We analyze the effect of $w_{r}$ by showing the embedding performance on the validation set of two different weight settings (uniform and tuned) on the Geology network. The results are shown in Figure 5. In the uniform setting, all relations have the same weight (i.e., $[w_{cb},w_{sa},w_{sv},w_{cr},w_{ccb}]=[1,1,1,1,1]$). We can find that as the learning step increases, cb, cr, and ccb embeddings quickly reach the best performance and then overfit, while sa and sv embeddings are still underfitting. This motivates us to increase the weight for sa and sv since their learning speeds are slow. We can find that when we use the tuned weight set ($[w_{cb},w_{sa},w_{sv},w_{cr},w_{ccb}]=[1,2,2,1,1]$), the learning speeds of different relations are more balanced, leading to a better general multiplex encoder.
@@ 6 Related Work
6.1 Learning Embeddings on Multiplex Networks
@@
Multiplex networks are also referred to as multi-view networks Qu et al. (2017); Shi et al. (2018) and multi-dimensional networks Berlingerio et al. (2013); Ma et al. (2018) in the literature. Multiplex networks consist of multiple relation/edge types among a set of single-type nodes. They can be seen as special cases of heterogeneous networks Dong et al. (2020); Sun et al. (2011) where nodes only have one type. The first category of existing works proposes to use one embedding vector to represent each node Dong et al. (2017); Hu et al. (2020); Wang et al. (2019); Zhang et al. (2019). HetGNN Zhang et al. (2019) introduces a graph neural network that utilizes heterogeneous neighbor aggregation to capture the diverse relation semantics. HAN Wang et al. (2019) proposes two-level attention including node-level and semantic-level to conduct relation semantics encoding. HGT Hu et al. (2020) introduces a more complex heterogeneous graph Transformer architecture to conduct node and edge semantics joint encoding. However, they all have an assumption that one embedding is enough to capture the semantics of all relations, which does not quite hold as shown in Section 3. The second category of works tries to develop multiplex embeddings for each node Jing et al. (2021); Park et al. (2020); Qu et al. (2017); Zhang et al. (2018), with one for each relation. MNE Zhang et al. (2018) proposes a random walk-based multiplex embedding learning method. MVE Qu et al. (2017) introduces a collaboration framework to capture the individual view network semantics and integrate them for the robust representations. DMGI Park et al. (2020) proposes a deep infomax framework to conduct multiplex embedding learning with graph neural networks. HDMI Jing et al. (2021) further enhances DMGI by introducing high-order mutual information with a fusion module. However, those works emphasize more on network structure encoding, while do not take the rich textual information associated with nodes into consideration.
6.2 Text Embeddings
@@
Text embeddings Brown et al. (2020); Cohan et al. (2020); Harris (1954); Le and Mikolov (2014); Ostendorff et al. (2022); Reimers and Gurevych (2019) effectively capture the textual semantic similarity between text units (i.e., sentences and documents) via distributed representation learning. Early work Harris (1954) proposes the bag-of-words vector space model, representing a text unit as a multiset of its words, disregarding grammar and even word order. Paragraph Vector Le and Mikolov (2014) is then introduced to capture the text semantics rather than just word appearance by representing each document with a dense vector trained to predict words in the document. As large-scale pretrained language models Beltagy et al. (2019); Brown et al. (2020); Devlin et al. (2019); Liu et al. (2019b) are proposed, Sentence-BERT Reimers and Gurevych (2019) further finetunes BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019b) by using a siamese and triplet network structure to derive semantically meaningful text embeddings which can measure text proximity with the cosine similarity between embeddings. OpenAI-ada-002 is recently developed as a powerful text embedding model based on the GPT series of large language models Brown et al. (2020). Specific to the academic domain, SPECTER Cohan et al. (2020) finetunes SciBERT by introducing positive and negative paper pairs, while SciNCL Ostendorff et al. (2022) further enhances SPECTER by developing a nearest neighbor sampling strategy. However, all the existing works presume that one embedding can capture the general semantics for each text unit and does not take the diverse text relation semantics into consideration.
@@ 7 Conclusions
We discuss the limitations of our work in Appendix A.1. In this work, we tackle the problem of representation learning on multiplex text-rich networks. To this end, we introduce Metern, learning multiplex text representations with only one text encoder. Metern introduces relation prior tokens to capture the relation-specific signals and one text encoder to model the shared knowledge across relations. We conduct experiments on nine downstream tasks and five networks from two domains, where Metern outperforms baselines significantly and consistently. Interesting future directions include: (1) exploring other more advanced graph-empowered text encoders for learning multiplex embeddings, and (2) applying the framework to more network-related tasks such as network generation conditioned on various relations.
@@ A.1 Limitations
In this work, we mainly focus on learning multiplex text/node representations on text-rich networks and solving downstream tasks (e.g., classification, regression, matching) with the learned embeddings. Because of the limited computational budgets, our PLM text encoder (bert-base-uncased) is medium-scale. In the future, we will explore applying a similar multiplex representation learning philosophy to large-scale language models. Other interesting future directions include designing more advanced graph-empowered language models to learn the multiplex embeddings and adopting the model into more real-world applications such as network generation.
@@ A.2 Ethical Considerations
PLMs have been shown to be highly effective in encoding contextualized semantics and understanding documents, as evidenced by several studies Devlin et al. [2019], Liu et al. [2019b], Clark et al. [2020]. However, some researchers have pointed out certain limitations associated with these models, including the presence of social bias Liang et al. [2021] and the propagation of misinformation Abid et al. [2021]. In our work, we focus on utilizing the relation signals between texts from the multiplex text-rich network structure to facilitate the understanding of the semantics of the texts, which we believe could help to address issues related to bias and misinformation.
@@ A.3 Datasets
The statistics of the five datasets can be found in Table 7. In academic networks, nodes correspond to papers and there are five types of relations between papers: “cited-by” (cb), “same-author” (sa), “same-venue” (sv), “co-reference” (cr), and “co-cited-by” (ccb); while in e-commerce networks, nodes are items and there are four types of relations between items: “co-purchased” (cop), “co-viewed” (cov), “bought-together” (bt), and “co-brand” (cob).
@@ A.4 Distribution Shift between Different Relations
$\text{Jac}(r_{k},r_{l})=\frac{|P_{r_{k}}(e_{ij})\cap P_{r_{l}}(e_{ij})|}{|P_{r_{k}}(e_{ij})\cup P_{r_{l}}(e_{ij})|}$ (11)
Since the whole networks are too large to calculate the Jaccard score, we randomly sample a sub-network from each network that contains 10,000 nodes and calculate Eq.(11). The results on Geology, Mathematics, Clothes, Home and Sports networks can be found in Figure 6. If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be nearly one, which is not the case in Figure 6.
More empirical experiments on the learned embedding distribution shift between relations in Mathematics, Clothes, Home, and Sports networks can be found in Figure 7. We finetune BERT [^4] We use the bert-base-uncased checkpoint. Devlin et al. [2019] to generate embeddings on one source relation distribution $P_{r_{k}}(e_{ij}|v_{i},v_{j})$ (row) and test the embeddings on the same or another target relation distribution $P_{r_{l}}(e_{ij}|v_{i},v_{j})$ (column). If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be nearly the same, which is not the case in Figure 7.
@@
Figure 6: Raw data distribution shift between different relations on Geology, Mathematic, Clothes, Home, and Sports network. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively. Each entry is the Jaccard score between the corresponding two relation distributions.
(a) Mathematics
(b) Clothes
(c) Home
(d) Sports
Figure 7: Distribution shift between different relations on Mathematic, Clothes, Home, and Sports network. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively. Each entry is the PREC@1 of BERT fine-tuned on the corresponding source relation distribution and tested on the corresponding target relation distribution.

### v2

@@ 5.4 Direct Inference with an Evident Source Relation
We conduct experiments on venue recommendation (Venue-Rec) and author identification (Author-Idf) in academic graphs and brand prediction (Brand-Pred) in e-commerce graphs to verify the effectiveness of Metag’s direct inference ability with an evident source relation. All three tasks are formulated as matching tasks where we use PREC@1 as the metric. For Venue-Rec, Author-Idf, and Brand-Pred, we use the sv, sa, and cob representation, respectively.
The results are shown in Table 3. From the results, we can find that: Metag outperforms all the baseline methods significantly and consistently on all the tasks (except Venue-Rec on Mathematics), which demonstrates that the multiplex embeddings generated by Metag are of good quality and can be directly applied to downstream tasks. Detailed information about the tasks and experimental settings can be found in Appendix A.5.2.
Table 3: Direct inference with an evident source relation (no task-specific training) on academic graphs and e-commerce graphs.
@@
Table 4: Learning to select source relations on academic graphs. ($\uparrow$): the greater the score is, the better the model is, and ($\downarrow$) otherwise.
@@
5.5 Learning to Select Source Relations
We perform experiments on paper recommendation, paper classification, citation prediction, and year prediction on academic graphs, and item classification and price prediction on e-commerce graphs to test if Metag can learn to select source relations. We use PREC@1 as the metric for paper recommendation, Macro-F1 as the metric for paper classification and item classification, and RMSE as the metric for citation prediction, year prediction, and price prediction.
Table 5: Learning to select source relations on academic graphs. ($\uparrow$) means the greater the score is, the better the model is, and ($\downarrow$) otherwise.
@@ 5.5 Learning to Select Source Relations
Attribute Vanilla FT MTDNN Metag
@@
The results on academic graphs and e-commerce graphs are shown in Table 4 and Table 6, respectively. We also show the learned source relation weight for citation prediction in Figure 3. From the results, we can find that: 1) Metag outperforms all the baseline methods significantly and consistently on all the tasks (except citation prediction on Geology and item classification on Sports). 2) Metag can learn to select different relations for different downstream tasks, i.e.,same-author” and “same-venue” for citation prediction (authors and venues can largely determine the citation of the paper). Detailed information about the tasks and experimental settings can be found in Appendix A.5.3. More result analysis can be found in Appendix A.7.
Figure 3: The learned source relation weights for citation prediction and paper recommendation on the Geology graph.
5.6 Efficiency Analysis
We run experiments on Geology to study the time complexity and memory complexity of Metag, comparing with Vanilla FT and MTDNN. All compared models are trained for 40 epochs on four NVIDIA A6000 GPU devices with a total training batch size set as 512. We show the result in Table 6. From the result, we can find that the time complexity and memory complexity of training Metag are on par with those of training Vanilla FT and MTDNN.
@@ 5.7 Multiplex Embedding Visualization
We visualize the multiplex node embeddings ${\bm{h}}_{v|r}$ learned by Metag with t-SNE (Van der Maaten & Hinton, 2008). The results on Geology are shown in Figure 4. We randomly select one center node from the graph, together with its neighboring nodes of the “cited-by” (cb), “same-author” (sa), and “same-venue” (sv) relations. Then, we utilize Metag to encode all the nodes (including the center node and all neighboring nodes) in different relation representation spaces (i.e., to obtain ${\bm{h}}_{v|r_{cb}}$, ${\bm{h}}_{v|r_{sa}}$, and ${\bm{h}}_{v|r_{sv}}$). In the figure, neighboring nodes of different relations are marked in different colors and shapes. From Figure 4, we can find that, in the output embedding space corresponding to one relation (e.g., the ${\bm{h}}_{v|r_{cb}}$ latent space corresponding to “cited-by” in Figure 4(a)), the center node (the blue circle)’s neighboring nodes of the corresponding relation (e.g., “cited-by” relation neighbors, marked as orange rectangles) are close to the center node. This demonstrates that Metag can capture the semantics of different relations and represent them in different latent subspaces.
@@
Figure 4: Multiplex embedding visualization on the Geology graph. cb, sa, and sv represent the “cited-by”, “same-author”, and “same-venue” relation respectively.
@@ 5.8 Analysis of Relation Weight $w_{r}$
In Metag, different relations sometimes have different learning speeds, and $w_{r}$ in Eq.(8) can be tuned to balance the learning speed of different relations. We analyze the effect of $w_{r}$ by showing the embedding performance on the validation set of two different weight settings (uniform and tuned) on the Geology graph. The results are shown in Figure 5. In the uniform setting, all relations have the same weight (i.e., $[w_{cb},w_{sa},w_{sv},w_{cr},w_{ccb}]=[1,1,1,1,1]$). We can find that as the learning step increases, cb, cr, and ccb embeddings quickly reach the best performance and then overfit, while sa and sv embeddings are still underfitting. This motivates us to increase the weight for sa and sv since their learning speeds are slow. We can find that when we use the tuned weight set ($[w_{cb},w_{sa},w_{sv},w_{cr},w_{ccb}]=[1,2,2,1,1]$), the learning speeds of different relations are more balanced, leading to a generally better multiplex encoder.
@@ 6 Related Work
6.1 Learning Embeddings on Multiplex Graphs
@@
Multiplex graphs are also referred to as multi-view graphs (Qu et al., 2017; Shi et al., 2018) and multi-dimensional graphs (Berlingerio et al., 2013; Ma et al., 2018) in the literature. Multiplex graphs consist of multiple relation/edge types among a set of single-typed nodes. They can be viewed as special cases of heterogeneous graphs (Dong et al., 2020; Sun et al., 2011) where nodes only have one type. The first category of existing works propose to use one embedding vector to represent each node (Dong et al., 2017; Hu et al., 2020; Wang et al., 2019; Zhang et al., 2019). For example, HetGNN (Zhang et al., 2019) introduces a graph neural network that utilizes heterogeneous neighbor aggregation to capture the diverse relation semantics; HAN (Wang et al., 2019) proposes two-level attention including node-level and semantic-level to conduct relation semantics encoding; HGT (Hu et al., 2020) introduces a more complex heterogeneous graph Transformer architecture to conduct node and edge semantics joint encoding. However, they all have an assumption that one embedding is enough to capture the semantics of all relations, which does not quite hold as shown in Section 3. The second category of works try to develop multiplex embeddings for each node (Jing et al., 2021; Park et al., 2020; Qu et al., 2017; Zhang et al., 2018), with one for each relation. For instance, MNE (Zhang et al., 2018) proposes a random walk-based multiplex embedding learning method; MVE (Qu et al., 2017) introduces a collaboration framework to capture the individual view graph semantics and integrate them for the robust representations; DMGI (Park et al., 2020) proposes a deep infomax framework to conduct multiplex embedding learning with graph neural networks; HDMI (Jing et al., 2021) further enhances DMGI by introducing high-order mutual information with a fusion module. However, those works emphasize more on graph structure encoding, while do not take the rich textual information associated with nodes into consideration.
6.2 Text Representations
@@
Text embeddings (Brown et al., 2020; Cohan et al., 2020; Harris, 1954; Le & Mikolov, 2014; Ostendorff et al., 2022; Reimers & Gurevych, 2019) effectively capture the textual semantic similarity between text units (i.e., sentences and documents) via distributed representation learning. Early work (Harris, 1954) proposes the bag-of-words vector space model, representing a text unit as a multiset of its words, disregarding grammar and even word order. Paragraph Vector (Le & Mikolov, 2014) is then introduced to capture the text semantics rather than just word appearance by representing each document with a dense vector trained to predict words in the document. As large-scale pretrained language models (Beltagy et al., 2019; Brown et al., 2020; Devlin et al., 2019; Liu et al., 2019b) are proposed, Sentence-BERT (Reimers & Gurevych, 2019) further finetunes BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019b) by using a siamese and triplet network structure to derive semantically meaningful text embeddings which can measure text proximity with the cosine similarity between embeddings. OpenAI-ada-002 is recently developed as a powerful text embedding model based on the GPT series of large language models (Brown et al., 2020). Specific to the academic domain, SPECTER (Cohan et al., 2020) finetunes SciBERT by introducing positive and negative paper pairs, while SciNCL (Ostendorff et al., 2022) further enhances SPECTER by developing a nearest neighbor sampling strategy. However, all the existing works presume that one embedding can capture the general semantics for each text unit and do not take the diverse text relation semantics into consideration.
@@ 7 Conclusions
In this work, we tackle the problem of representation learning on multiplex text-attributed graphs. To this end, we introduce the Metag framework to learn multiplex text representations with only one text encoder. Metag introduces relation prior tokens to capture the relation-specific signals and one text encoder to model the shared knowledge across relations. We conduct experiments on nine downstream tasks and five graphs from two domains, where Metag outperforms baselines significantly and consistently. Interesting future directions include: (1) exploring more advanced graph-empowered text encoders for learning multiplex embeddings, and (2) applying the framework to more graph-related tasks such as multiplex graph generation.
Impact Statements
In this work, we mainly focus on learning multiplex text/node representations on text-attributed graphs and solving downstream tasks (e.g., classification, regression, matching) with the learned embeddings. Because of the limited computational budgets, our PLM text encoder (bert-base-uncased) is medium-scale. In the future, we will explore applying a similar multiplex representation learning philosophy to large-scale language models. Other interesting future directions include designing more advanced graph-empowered language models to learn the multiplex embeddings and adopting the model into more real-world applications such as graph generation.
PLMs have been shown to be highly effective in encoding contextualized semantics and understanding documents, as evidenced by several studies (Devlin et al., 2019; Liu et al., 2019b; Clark et al., 2020). However, some researchers have pointed out certain limitations associated with these models, including the presence of social bias (Liang et al., 2021) and the propagation of misinformation (Abid et al., 2021). In our work, we focus on utilizing the relation signals between texts from the multiplex text-attributed graph structure to facilitate the understanding of the semantics of the texts, which we believe could help to address issues related to bias and misinformation.
@@ A.1 Limitations
In this work, we mainly focus on learning multiplex text/node representations on text-attributed graphs and solving downstream tasks (e.g., classification, regression, matching) with the learned embeddings. Because of the limited computational budgets, our PLM text encoder (bert-base-uncased) is medium-scale. In the future, we will explore applying a similar multiplex representation learning philosophy to large-scale language models. Other interesting future directions include designing more advanced graph-empowered language models to learn the multiplex embeddings and adopting the model into more real-world applications such as graph generation.
@@ A.2 Ethical Considerations
PLMs have been shown to be highly effective in encoding contextualized semantics and understanding documents, as evidenced by several studies (Devlin et al., 2019; Liu et al., 2019b; Clark et al., 2020). However, some researchers have pointed out certain limitations associated with these models, including the presence of social bias (Liang et al., 2021) and the propagation of misinformation (Abid et al., 2021). In our work, we focus on utilizing the relation signals between texts from the multiplex text-attributed graph structure to facilitate the understanding of the semantics of the texts, which we believe could help to address issues related to bias and misinformation.
@@ A.3 Datasets
The statistics of the five datasets can be found in Table 7. In academic graphs, nodes correspond to papers and there are five types of relations between papers: “cited-by” (cb), “same-author” (sa), “same-venue” (sv), “co-reference” (cr), and “co-cited-by” (ccb); while in e-commerce graphs, nodes are items and there are four types of relations between items: “co-purchased” (cop), “co-viewed” (cov), “bought-together” (bt), and “co-brand” (cob).
@@ A.4 Distribution Shift between Different Relations
$\text{Jac}(r_{k},r_{l})=\frac{|P_{r_{k}}(e_{ij})\cap P_{r_{l}}(e_{ij})|}{|P_{r_{k}}(e_{ij})\cup P_{r_{l}}(e_{ij})|}.$ (13)
Since the whole graphs are too large to calculate the Jaccard score, we randomly sample a sub-graph from each graph that contains 10,000 nodes and calculate Eq.(13). The results on Geology, Mathematics, Clothes, Home, and Sports graphs can be found in Figure 6. If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be nearly one, which is not the case in Figure 6.
More empirical experiments on the learned embedding distribution shift between relations in Mathematics, Clothes, Home, and Sports graphs can be found in Figure 7. We finetune BERT [^4] We use the bert-base-uncased checkpoint. (Devlin et al., 2019) to generate embeddings on one source relation distribution $P_{r_{k}}(e_{ij}|v_{i},v_{j})$ (row) and test the embeddings on the same or another target relation distribution $P_{r_{l}}(e_{ij}|v_{i},v_{j})$ (column). If the assumption of analogous distributions (i.e., $P_{r_{k}}(e_{ij}|v_{i},v_{j})\approx P_{r_{l}}(e_{ij}|v_{i},v_{j})$) holds, the values in each cell should be nearly the same, which is not the case in Figure 7.
@@
Figure 6: Raw data distribution shift between different relations on Geology, Mathematic, Clothes, Home, and Sports graphs. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively. Each entry is the Jaccard score between the corresponding two relation distributions.
(a) Geology
(b) Mathematics
(c) Clothes
(d) Home
(e) Sports
Figure 7: Distribution shift between different relations on Geology, Mathematic, Clothes, Home, and Sports graphs. cb, sa, sv, cr, and ccb represent “cited-by”, “same-author”, “same-venue”, “co-reference”, and “co-cited-by” relation respectively. cop, cov, bt, and cob represent “co-purchased”, “co-viewed”, “bought-together”, and “co-brand” relation respectively. Each entry is the PREC@1 of BERT fine-tuned on the corresponding source relation distribution and tested on the corresponding target relation distribution.